# Will Humanity Be Rendered Obsolete by AI?

### v1

@@ Abstract
This article analyzes the existential risks artificial intelligence (AI) poses to humanity, tracing the trajectory from current AI to ultraintelligence. Drawing on Irving J. Good and Nick Bostrom’s theoretical work, plus recent publications (AI 2027; If Anyone Builds It, Everyone Dies), it explores AGI and superintelligence. Considering machines’ exponentially growing cognitive power and hypothetical IQs, it addresses the ethical and existential implications of an intelligence vastly exceeding humanity’s, fundamentally alien. Human extinction may result not from malice, but from uncontrollable, indifferent cognitive superiority.
@@ Conclusion
But if we consider the concept of intelligence explosion and exponential dynamics, everything depends on when the AI becomes incomparably smarter than us. At an IQ of 200, humans remain competitors. At 1000, they become invisible. Future work should explore mechanisms for Value Lock-in (Bostrom, 2014; Soares and Fallenstein, 2017; Christian, 2020; Harari, 2024) capable of resisting an intelligence explosion, as well as the possibilities of human-machine cognitive symbiosis as an alternative to algorithmic supremacy.
@@
Humanity thus faces a dilemma: voluntarily renounce a part of its cognitive power by deliberately limiting its own technological advances to maintain control, or continue to delegate to intelligences that will soon no longer need it. The major danger, therefore, does not lie in the machine itself but in our inability to define limits to what we still consider progress. Will humanity have to curb progress to avoid losing control?
@@
Humans assess the situation using their own intelligence. They are simply unfamiliar with intelligences not originating from their own minds, and are fundamentally wary of them. For centuries, they have measured intelligence with their own yardstick, using tools like IQ, overlooking other aspects of our own intelligence and even other forms of intelligence. It is possible to place certain animals and plants at a higher level than humans in terms of living being intelligence (animate intelligence, Lent, 2021). We have ignored that domestic animals, elephants, and dolphins, among others, dream, and that forests have a memory. In many areas, some animals, like orangutans, surpass humans in intelligence, particularly for forest survival. Even among humans, it is well established that IQ does not encompass all forms of intelligence, as defined by Gardner (1993), for example.
Following the intelligence explosion, the fact that humans no longer know what this concept will mean invalidates Minsky’s observation (1968) that AI is only intelligent insofar as it tackles problems that humans solve thanks to intelligence; he defined AI according to the metric of human intelligence. This observation, linking AI to human intelligence, leads us to raise questions about the true nature of renewing intelligence. As AI evolves, it becomes less artificial (in the sense that it depends on human conceptions) and more foreign (see Harari’s discussion on this subject, 2024, particularly on page 320). The last act of human consciousness could then consist in contemplating, with a mixture of astonishment and admiration, the emergence of its successor, perhaps congratulating it for its success.

### v2

@@ Abstract
This article analyzes the existential risks Artificial Intelligence (AI) poses to humanity, tracing the trajectory from current AI to ultraintelligence. Drawing on Irving J. Good and Nick Bostrom’s theoretical work, plus recent publications (AI 2027; If Anyone Builds It, Everyone Dies), it explores Artificial General Intelligence (AGI) and superintelligence. Considering machines’ exponentially growing cognitive power and hypothetical IQs, it addresses the ethical and existential implications of an intelligence vastly exceeding humanity’s, fundamentally alien. Human extinction may result not from malice, but from uncontrollable, indifferent cognitive superiority.
@@ Introduction
Zuboff (2019) already spoke of the “Big Other”, a distributed and largely automatic apparatus that replaces the totalitarian logic of Orwell’s “Big Brother” with a new form of power, an “instrumentarian” power, that aims not at overt domination but at the automated shaping of human behavior (see pp. 367-377).
@@ Conclusion
But if we consider the concept of intelligence explosion and exponential dynamics, everything depends on when the AI becomes incomparably smarter than us. At an IQ of 200, humans remain competitors. At 1000, we become invisible. Future work should explore mechanisms for Value Lock-in (Bostrom, 2014; Soares and Fallenstein, 2017; Christian, 2020; Harari, 2024) capable of resisting an intelligence explosion, as well as the possibilities of human-machine cognitive symbiosis as an alternative to algorithmic supremacy.
@@
Humanity thus faces a dilemma: voluntarily renounce a part of its cognitive power by deliberately limiting its own technological advances to maintain control, or continue to delegate to intelligences that will soon no longer need it. The major danger, therefore, does not lie in the machine itself but in our inability to define limits to what we still consider progress. Will we have to curb progress to avoid losing control?
@@
Humans assess the situation using their own intelligence. They are simply unfamiliar with intelligences not originating from their own minds, and are fundamentally wary of them. For centuries, they have measured intelligence with their own yardstick, using tools like IQ, overlooking other aspects of their own intelligence and even other forms of intelligence. Even among humans, it is well established that IQ does not encompass all forms of intelligence, as defined by Gardner (1993), for example.
Following the intelligence explosion, the fact that humans no longer know what this concept will mean invalidates Minsky’s observation (1968) that AI is only intelligent insofar as it tackles problems that humans solve thanks to intelligence; he defined AI according to the metric of human intelligence. This observation, linking AI to human intelligence, leads us to raise questions about the true nature of renewing intelligence. As AI evolves, it becomes less artificial (in the sense that it depends on human conceptions) and more foreign (see Harari’s discussion on this subject, 2024, particularly on page 320). Human consciousness could then involve observing, with interest, the emergence of this new intelligence.